The electron in a hydrogen atom can undergo a transition from n=1 to n=6, absorbing a photon with a wavelength of 94 nm.

How much energy must be absorbed for this transition to occur?

How does this transition show that the energy of a photon is quantized?

How does this absorption begin to approximate the ionization energy of hydrogen?
Answer #1

(a): Energy absorbed, E = h*c / = 6.626*10-34 J.s* 3*108 m.s-1/ 9.4*10-8 m = 2.115*10-18 J

(b): Energy radiated during transition from n=1 to n=6

E = - 2.18*10-18J*(1/62 - 1/1)

=> E = - 2.18*10-18J* (-0.9722)

=> E = 2.119*10-18J

Since energy radiated and energy absorbed are approximately same, energy is quantized.

(c): Ionization energy of H is 2.18*10-18J.

Since the above ionization energy is approximately same as 2.119*10-18J, this absorption begin to approximate the ionization energy of hydrogen.